# Generating and Customizing Robotic Arm Trajectories using Neural Networks

### v1

@@ 2.1 Formal definition of the problem
Our method requires measuring the start and end poses of the robotic arm (given by joint angles), the shape of the trajectory, and the number of steps to control the motors. The start poses $\boldsymbol{p}^{\rm s}=(\theta^{\rm s}_{1},\theta^{\rm s}_{2},\dots,\theta^{\rm s}_{m})$, and the end poses $\boldsymbol{p}^{\rm e}=(\theta^{\rm e}_{1},\theta^{\rm e}_{2},\dots,\theta^{\rm e}_{m})$ are tuples of angles; in our particular case, we use $m=7$ degrees of freedom. From these poses, we calculate the goal start point $\boldsymbol{P}^{\rm s}=(x_{\rm s},y_{\rm s},z_{\rm s})$ and the goal end point $\boldsymbol{P}^{\rm e}=(x_{\rm e},y_{\rm e},z_{\rm e})$ by FK described below. We translate the trajectory of the desired shape with these points and divide it into the number of $n$ segments. This division must be accurate enough to assume constant angular velocities for individual motors within one segment. However, the duration of these segments must correspond to our ability to poll the motors via the bus they are connected to. We have to choose $n$ to balance the two requirements.
@@
In addition, we can derive $n$ goal vectors $\boldsymbol{v}^{\rm g}_{i}$ from the $n+1$ goal points as $\boldsymbol{v}^{\rm g}_{i}=\boldsymbol{P}^{\rm g}_{i+1}-\boldsymbol{P}^{\rm g}_{i}$ for $i=0,1,\dots,n-1$. All vectors are identical when the trajectory shape is a line; in this case, $\boldsymbol{v}_{i}=(\boldsymbol{P}^{\rm e}-\boldsymbol{P}^{\rm s})/n$. The method provides us with $n+1$ poses of the robot arm (joint angles) $\{(\theta^{i}_{1},\theta^{i}_{2},\dots\theta^{i}_{m}\mid i=0,1,\dots,n\}$. From them we can easily calculate the corresponding angular velocities that move all the motors during the whole segment, taking $T/n$ ms, where $T$ is the duration of the whole movement in milliseconds.
@@ 2.2 Neural network architecture for forward kinematics
Similarly, when using $3\times 3$ parts of the rotational matrices, we can project the vector $(0,0,1)$ to get joint orientations, including the roll, pitch and yaw of the end effector.
@@
Unlike the traditional approach, we never evaluate the product of all matrices $M_{k-1}\times\dots\times M_{1}\times M_{0}$ (see Eq. 2). We always multiply the initial vector ${(0,0,0,1)}^{\top}$ by matrices one by one: $(M_{k-1}\times\dots\times(M_{1}\times(M_{0}\times{(0,0,0,1)}^{\top}))\dots)$. The most important deviation is that the input to this module is not the initial vector but the joint angles $\theta_{j}$. We apply cosine and sine on them, update the matrices, and output the calculated position and orientation of the end effector. So, though the matrices linearly process vectors, this module is made to be a very nonlinear module of a neural network. Neither parameter of this module is trainable. However, we can run it not only in the forward mode that calculates point and orientation from angles but also in the backward mode that calculates angles’ gradients upon the deviation of the point and orientation from their goal values. Also, it is essential to implement this module to receive not only a single input but also a batch, as is usual for neural networks.
@@ 2.3 Neural network architecture for inverse kinematics
With the module mentioned above, we can turn the IK problem (finding joint angles for a given 3D point) into the training of a neural network consisting of two parts: a generator of angles and the FK module. The generator provides proposals for angles, and the FK turns them into the corresponding 3D point. Then, we aim to adjust the generator parameters to decrease the distance between the predicted point and the target.
Since we seek to train a neural network, we must formally define a dataset, while we know only one desired output (the given point). We can treat this requirement by assigning a formal input to the generator, always $1$. We process it by a linear layer of $m$ neurons, each having one weight and no bias, whose weights code the angles.
The primary design challenge for the generator is operationalise the weights training while preserving the angle ranges. We need to ensure that the angles are within their range, can reach the extreme value, and break away from it as required during training. Clipping the angle values is not a solution due to zero gradients at the extremes. Therefore, first, we generate angles in the form of logits $z\in(-\infty,\infty)$, and then we apply a sigmoid and turn them to degrees, considering their range:
$\theta_{j}=\theta_{j}^{\rm min}+\sigma(z)(\theta_{j}^{\rm max}-\theta_{j}^{\rm min})$ (13)
If we create a dataset with a single sample of input $1$ and output $x,y,z$, we could train the model until the weights in the linear layer correspond to the proper logits that provide the desired output point. Then, we can use the above formula to get the corresponding values in degrees that form the required pose of the robotic arm.
@@ 2.4 Neural network architecture for trajectory generation
  • the mean distance between points $\boldsymbol{P}_{i}$ and goals $\boldsymbol{P}^{\rm g}_{i}$:
@@
  • negative mean of cosines of angles between vectors $\boldsymbol{v}_{i}$ corresponding to the calculated points $\boldsymbol{P}_{i}$, $\boldsymbol{v}_{i}=\boldsymbol{P}_{i+1}-\boldsymbol{P}_{i}$ and the goal vectors $\boldsymbol{v}^{\rm g}_{i}=\boldsymbol{P}^{\rm g}_{i+1}-\boldsymbol{P}^{\rm g}_{i}$: $\mathcal{L}_{1}=1-\frac{1}{n}\sum\limits_{i=0}^{n-1}{\frac{\boldsymbol{v}_{i}^{\top}\boldsymbol{v}^{\rm g}_{i}}{\lVert\boldsymbol{v}_{i}\rVert\lVert\boldsymbol{v}^{\rm g}_{i}\rVert}}$
@@
Generating the whole trajectory with one training process requires slightly modifying the dataset for the single IK problem. The dataset again contains a single sample. Its input is a vector of $n+1$ ones, and the desired output is the vector of the $n+1$ goal points. We use Adam optimizer with learning rate $0.1$. The initial setup of weights representing logits of angles follows analogical mixtures of the start and end poses as we mix the goal points from the start and endpoint. This training gradually achieves the required shape of the trajectory and improves its fluency. The training transforms an initial trajectory designed in the joint space of the robot to the trajectory of the required shape in the Cartesian space. Therefore, this method retrieves the 3D coordinates that the robot’s end effector should assume to follow the required trajectory. We can stop training when the angles do not change enough. We can also stop when the generated trajectory has a sufficient shape. The duration of training varies from seconds to minutes; it depends mainly on the end point and the availability of the sought trajectory.
@@ 3 Results
Following the objectives of the experiment [18], we generated seven trajectories to touch seven selected points on the touchscreen. They started from the same position with the raised arm and provided pointing by the forefinger to the point of touch throughout the movement, except in the starting phase. We recorded the start pose and seven touching poses; the NICO robot enables switching torque off and on and reading the current angles anytime, so managing this is quite comfortable. Training generated the trajectories successfully, as shown in Figure 2. The number of iterations is provided in Table 1. The NICO robot has only one degree of freedom at the elbow. As a result, it isn’t easy to reach some spots with its arm, namely spots close to the robot; this may be why closer points require a much higher number of iterations.
Table 1: Duration and accuracy of trajectories generation
@@
Table 2: Duration and accuracy of trajectories generation

### v2

@@ 2.1 Formal definition of the problem
Our method requires measuring the start and end poses of the robotic arm (given by joint angles), the shape of the trajectory, and the number of steps to control the motors. The start pose $\boldsymbol{p}^{\rm s}=(\theta^{\rm s}_{1},\theta^{\rm s}_{2},\dots,\theta^{\rm s}_{m})$ and the end pose $\boldsymbol{p}^{\rm e}=(\theta^{\rm e}_{1},\theta^{\rm e}_{2},\dots,\theta^{\rm e}_{m})$ are tuples of angles; in our particular case, we use $m=7$ degrees of freedom. From these poses, we calculate the goal start point $\boldsymbol{P}^{\rm s}=(x_{\rm s},y_{\rm s},z_{\rm s})$ and the goal end point $\boldsymbol{P}^{\rm e}=(x_{\rm e},y_{\rm e},z_{\rm e})$ by FK described below. We translate the trajectory of the desired shape with these points and divide it into $n$ segments. This division must be accurate enough to assume constant angular velocities for individual motors within one segment. However, the duration of these segments must correspond to our ability to poll the motors via the bus they are connected to. We have to choose $n$ to balance the two requirements.
@@
In addition, we can derive $n$ goal vectors $\boldsymbol{v}^{\rm g}_{i}$ from the $n+1$ goal points as $\boldsymbol{v}^{\rm g}_{i}=\boldsymbol{P}^{\rm g}_{i+1}-\boldsymbol{P}^{\rm g}_{i}$ for $i=0,1,\dots,n-1$. All vectors are identical when the trajectory shape is a line; in this case, $\boldsymbol{v}_{i}=(\boldsymbol{P}^{\rm e}-\boldsymbol{P}^{\rm s})/n$. The method provides us with $n+1$ poses of the robot arm (joint angles) $\{(\theta^{i}_{1},\theta^{i}_{2},\dots\theta^{i}_{m}\mid i=0,1,\dots,n\}$. From them, we can easily calculate the corresponding angular velocities that move all the motors during the whole segment, taking $T/n$ ms, where $T$ is the selected duration of the whole movement in milliseconds.
@@ 2.2 Neural network architecture for forward kinematics
Similarly, when using $3\times 3$ parts of the rotational matrices, we can project the vector $(0,0,1)$ to get joint orientations, including the roll, pitch, and yaw of the end effector.
@@
Unlike the traditional approach, we never evaluate the product of all matrices $M_{k-1}\times\dots\times M_{1}\times M_{0}$ (see Eq. 2). We always multiply the initial vector ${(0,0,0,1)}^{\top}$ by matrices one by one: $(M_{k-1}\times\dots\times(M_{1}\times(M_{0}\times{(0,0,0,1)}^{\top}))\dots)$. The most important deviation is that the input to this module is not the initial vector but the joint angles $\theta_{j}$. We apply cosine and sine on them, update the matrices, and output the calculated position and orientation of the end effector. So, though the matrices linearly process vectors, this module is made to be a very nonlinear module of a neural network. Neither parameter of this module is trainable (although it could be within another project). However, we can run it not only in the forward mode that calculates point and orientation from angles but also in the backward mode that calculates angles’ gradients upon the deviation of the point and orientation from their goal values. Also, it is essential to implement this module to receive not only a single input but also a batch, as is usual for neural networks.
@@ 2.3 Neural network architecture for inverse kinematics
With the module mentioned above, we can turn the IK problem (finding joint angles for a given 3D point and orientation) into training a neural network consisting of two parts: a generator of angles and the FK module. The generator provides proposals for angles, and the FK transforms them into the corresponding 3D point and orientation. Then, we aim to adjust the generator parameters to decrease the distance between the predicted values and the target.
Since we seek to train a neural network, we must formally define a dataset, while we know only one desired output (the given point and orientation). We can treat this requirement by assigning a formal input to the generator, always $1$. We process it by a linear layer of $m$ neurons, each having one weight and no bias, whose weights code the angles.
The primary design challenge for the generator is to operationalise the weights training while preserving the angle ranges. We need to ensure that the angles are within their range, can reach the extreme value, and break away from it as required during training. Clipping the angle values is not a solution due to zero gradients at the extremes. Therefore, first, we generate angles in the form of logits $z_{j}\in(-\infty,\infty)$, and then we apply a sigmoid and turn them to degrees, considering their range:
$\theta_{j}=\theta_{j}^{\rm min}+\sigma(z_{j})(\theta_{j}^{\rm max}-\theta_{j}^{\rm min})$ (13)
If we create a dataset with a single sample of input $1$ and output $x$, $y$, $z$, roll, pitch, and yaw, we could train the model until the weights in the linear layer correspond to the proper logits that provide the desired output. Then, we can use the above formula to get the corresponding values in degrees that form the required pose of the robotic arm.
@@ 2.4 Neural network architecture for trajectory generation
  • the mean distance between output points $\boldsymbol{P}_{i}$ and goal points $\boldsymbol{P}^{\rm g}_{i}$:
@@
  • negative mean of cosines of angles between output orientation vectors $\boldsymbol{v}_{i}$ and goal vectors $\boldsymbol{v}^{\rm g}_{i}$ (in our particular case of the linear shape $\boldsymbol{v}^{\rm g}_{i}=(\boldsymbol{P}^{\rm e}-\boldsymbol{P}^{\rm s})/n$): $\mathcal{L}_{1}=1-\frac{1}{n+1}\sum\limits_{i=0}^{n}{\frac{\boldsymbol{v}_{i}^{\top}\boldsymbol{v}^{\rm g}_{i}}{\lVert\boldsymbol{v}_{i}\rVert\lVert\boldsymbol{v}^{\rm g}_{i}\rVert}}$
@@
Generating the whole trajectory with one training process requires slightly modifying the dataset for the single IK problem. The dataset again contains a single sample. Its input is a vector of $n+1$ ones, and the desired output is the vector of the $n+1$ goal points and orientations. We use Adam optimizer with learning rate $0.1$. The initial setup of weights representing logits of angles follows analogical mixtures of the start and end poses as we mix the goal points from the start and endpoint. This training gradually achieves the required shape of the trajectory and improves its fluency. The training transforms an initial trajectory designed in the joint space of the robot to the trajectory of the required shape in the Cartesian space. Therefore, this method retrieves the 3D coordinates that the robot’s end effector should assume to follow the required trajectory. We can stop training when the angles do not change enough. We can also stop when the generated trajectory has a sufficient shape. The duration of training varies from seconds to minutes; it depends mainly on the end point and the availability of the sought trajectory.
@@ 3 Results
Following the objectives of the experiment [18], we generated seven trajectories to touch seven selected points on the touchscreen. They started from the same position with the raised arm and provided pointing by the forefinger to the point of touch throughout the movement, except in the starting phase. We recorded the start pose and seven touching poses. The NICO robot enables switching torque off and on and reading the current angles anytime, so managing this is quite comfortable. Training generated the trajectories successfully, as shown in Figure 2. The number of iterations is provided in Table 1. Since the NICO robot has only one degree of freedom at the elbow, it isn’t easy to reach some spots with its arm, namely spots close to the robot. It may be the reason why closer points require a much higher number of iterations.
Table 1: Duration and accuracy of trajectories generation (our method)
@@
Table 2: Accuracy of trajectories generation (the traditional method)